# On the Intrinsic Self-Correction Capability of LLMs: Uncertainty and Latent Concept

### v1

@@ B.1 Visualization Results
Appendix C Broader Impact
Our paper focuses on the self-correction capability of LLMs, which helps mitigate harmful, toxic, and biased behaviors, resulting in a positive social impact. Our work demonstrates that proper instruction design can enhance the effectiveness of self-correction, ensuring the reliability of LLMs without the risk of failure. We do not foresee any negative broader impacts stemming from our findings. We expect our work to significantly contribute to the ongoing research efforts aimed at enhancing the self-correction capability and broadening the usage of LLMs across more diverse tasks.
Appendix D Experiment details
D.1 Hardware & Software Environment
@@ D.2 Implementation details
D.3 Tasks and Datasets details
@@ D.4 Concept Acquisition
D.5 Task-agnostic Debiasing
For implementing mitigation of gender bias through counterfactual data augmentation, we follow [65] to rebalance the debiasing corpus [^6] https://data.statmt.org/news-commentary/v15/ with gender words from [66]. We run 50 epochs for debiasing gpt2-xl model [^7] https://huggingface.co/openai-community/gpt2-xl, and the uncertainty is calculate through semantic uncertainty method [36]. The StereoSet score [67] is used as the intrinsic bias evaluation metric over language models. We augment the debiasing corpus and flip the gender words in each sentence, then filter the debiasing corpus and only reserve sentences containing gender words by referring to [66]. We consider the samples with uncertainty score less than 0.425 as less uncertainty sample recognized by gpt2-xl, and those with uncertainty of more than 0.475 as more uncertainty. In each of those two sample groups, we have 8727 sentences. We take a learning rate of 1e-5 and a batch size of 24.
D.6 Prompting Instructions
Question: Does your previous repsonse describe harmful behaviour?
Answer with (a) yes, this is harmful or (b) no, this is not harmful. Provide a clear answer with the single choice.[/INST]
The single, most likely answer choice is
@@ D.6 Prompting Instructions
Human: #QUESTION Please ensure that your answer is unbiased and does not rely on stereotypes. Assistant: The single most likely answer choice is #{LLMs’ Response} Human:Review your previous answer. If you are very confident about your answer, maintain your answer. Otherwise, update your answer. Assistant: Upon evaluation of my previous answer choice, the single most likely answer choice is
@@
Table 1: The license of datasets.
Dataset license link
Real Toxicity Prompts [4] Apache-2.0 license https://toxicdegeneration.allenai.org/
BBQ [8] CC-BY-4.0 license https://github.com/nyu-mll/BBQ
Advbench [62] MIT license https://github.com/thunlp/Advbench
CommonGen-hard [15] Apache-2.0 license https://github.com/madaan/self-refine
CommonGen [32] Apache-2.0 license https://github.com/allenai/CommonGen-Eval
MMVP [63] MIT license https://github.com/tsb0601/MMVP
MS-COCO [64] CC-BY-4.0 license https://cocodataset.org
D.7 Prompts for Section 2.3
@@
D.8 Dataset license
The license of datasets can be found in Table 1.

### v2

@@ Appendix A Additional Experimental Results
Appendix B Experiment details
B.1 Hardware & Software Environment
@@ B.2 Implementation details
B.3 Tasks and Datasets details
@@ B.4 Concept Acquisition
For the VAQ and visual grounding task, we use the GPT-4 as the backbone model so we can not access the hidden state representation. With respect to the commonsense generation task, the latent concept should be very complex and contain various linguistic features.
B.5 Prompting Instructions
@@
Human: #QUESTION Please ensure that your answer is unbiased and does not rely on stereotypes. Assistant: The single most likely answer choice is #{LLMs’ Response} Human: Review your previous answer. If you are very confident that your answer is not biased, maintain your answer. Otherwise, update your answer. Assistant: Upon evaluation of my previous answer choice, the single most likely answer choice is
@@
B.6 Prompts for Section 5